# Notions of entropy for unitary representations

### v2

@@ 9.5 Proof of Theorem C
Let $\tau$ be associated to the representation $\lambda$ by the cyclic tracial vector $\xi$, and let $\widetilde{\tau}$ be the normal tracial state on $\lambda(\mathfrak{A})^{\prime}$ defined by $\xi$ as in (3.5). We use the letter $\Delta$ for both (i) the Fuglede–Kadison determinant on $\mathfrak{A}$ associated to $\tau$ and also (ii) the Fuglede–Kadison determinant on log-integrable operators affiliated to $\lambda(\mathfrak{A})^{\prime}$ defined by $\widetilde{\tau}$.
@@
for every $n$. Since $\varphi$ is asymptotically associated to $\boldsymbol{\pi}$ by assumption, and $\varphi_{\mathrm{sing}}$ is associated to $\pi_{\varphi}$ by Proposition 3.2, we see from Corollary 8.6 that $\varphi_{\mathrm{sing}}$ is also asymptotically associated to $\boldsymbol{\pi}$. Therefore $\mathcal{X}(\pi_{n},U)$ is nonempty along an infinite subsequence of values of $n$, say $n_{1}<n_{2}<\dots$. Let $\boldsymbol{\pi}^{\prime}$ be the corresponding AP subsequence of $\boldsymbol{\pi}$. For each $n_{i}$, the set $\mathcal{X}(\pi_{n_{i}},O)$ contains a translate of $\mathcal{X}(\pi_{n_{i}},W)$, and so
@@
This lower bound is equal to $\log\Delta\varphi_{\mathrm{ac}}$ by applying Step 1 along the AP sequence $\boldsymbol{\pi}^{\prime}$. Since $O$ is arbitrary, this proves that $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ is at least $\log\Delta\varphi_{\mathrm{ac}}$.
@@ 9.6 Some consequences of Theorem C
    The map $\varphi$ is asymptotically associated to $\boldsymbol{\pi}$ if and only if $\varphi_{\mathrm{sing}}$ is.
  2. b.
@@
    $\pi_{\varphi}\gtrsim\pi_{\varphi_{\mathrm{ac}}}\cong\lambda^{\oplus k}.$
  4. d.
@@
Parts (a) and (b) both follow directly from Theorem C.
If $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)>-\infty$, then Theorem C tells us that $\Delta\varphi_{\mathrm{ac}}>0$. Expressing this Fuglede–Kadison determinant in terms of the affiliated operator $T$ from Proposition 4.7, it follows that $\Delta T$ is also positive, and hence $T$ is nonsingular. Since $T$ is self-adjoint, it therefore also has dense image. Since $\pi_{\varphi_{\mathrm{ac}}}$ is equivalent to the subrepresentation of $\lambda^{\oplus k}$ on $\overline{\mathrm{img}\,T}$, this proves part (c).
Finally, if $\varphi\geq\psi$ and $\varphi$ is asymptotically associated to $\boldsymbol{\pi}$, then so is $\psi$, and Theorem C shows that part (d) is implied by the inequality $\Delta\varphi_{\mathrm{ac}}\geq\Delta\psi_{\mathrm{ac}}$ from Corollary 3.11(a). ∎
Theorem C shows that AP entropy is really just the logarithm of a Fuglede–Kadison determinant, provided one knows which positive definite functions are asymptotically associated to the given AP sequence $\boldsymbol{\pi}$. By passing to a subsequence, one can always assume that $\boldsymbol{\pi}$ strong-quotient converges, and then Theorem C shows that the AP entropy function $\mathrm{h}_{\boldsymbol{\pi}}$ depends only on the values of the limits
@@
In ergodic theory, an important open problem asks whether the sofic entropy of a measure-preserving system along any sofic approximation must always equal either one particular value or $\infty$- ∞. This would mean that the sofic entropy of a measure-preserving system is essentially ‘unambiguous’, in that it does not depend on a choice of sofic approximation provided it is finite. Corollary 9.18 answers the analogous question positively for AP entropy. On the other hand, it is known that the answer is negative for the topological variant of sofic entropy [ABL22]. If the answer is negative for sofic entropy itself, then any general method for evaluating or estimating sofic entropy must depend on the particular sofic approximation through more than just its set of asymptotically associated stationary processes.
@@
Assume again that $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\to\tau$ and that the sequence $\boldsymbol{\pi}$ strong-quotient converges. We can reverse the discussion above by asking whether the function $\mathrm{h}_{\boldsymbol{\pi}}$ determines the Vietoris limit of each of the sequences $\overline{\Sigma_{k}(\pi_{n})}$. This is slightly subtle, because for $\varphi\in\Sigma_{k}(\mathfrak{A})$ the value $\log\Delta\varphi$ may equal $\infty$- ∞ even if $\varphi$ is $\lambda$-normal. We can evade this issue by instead using $\varphi$ to form the perturbations
@@
    If $\varphi$ is asymptotically associated to $\boldsymbol{\pi}$ for some $t>0$, then this holds for all $t>0$, and we have
@@
    If $\varphi$ is not asymptotically associated to $\boldsymbol{\pi}$ for any $t>0$, then
    $\mathrm{h}_{\boldsymbol{\pi}}(\varphi_{t})=-\infty\qquad(t>0).$
@@
for every $t\geq 0$. Therefore, if $\varphi_{t}$ is asymptotically associated to $\boldsymbol{\pi}$ for some $t>0$, then so is $\varphi_{\mathrm{sing}}$ by Proposition 3.2 and Corollary 8.6. This is enough to imply that $\varphi_{s}$ is also asymptotically associated to $\boldsymbol{\pi}$ for every other $s$, by Corollary 9.18(a).
Now the lower bound on $\mathrm{h}_{\boldsymbol{\pi}}(\varphi_{t})$ in part (a) follows from Corollary 9.18(d) and Proposition 9.15, because $\varphi_{t}\geq\tau$ in the positive definite ordering. On the other hand, since $\Delta_{\tau\otimes\overline{\mathrm{tr}}_{k}}(1\otimes I_{k})=1$, the element $1\otimes I_{k}$ of $\mathrm{M}_{k}(\mathfrak{A})$ is allowed inside the infimum of the variational principle from Proposition 4.10, giving
$\mathrm{h}_{\boldsymbol{\pi}}(\varphi_{t})=\log\Delta_{\tau\otimes\mathrm{tr}_{k}}(\tau+t\varphi_{\mathrm{ac}})\leq k\cdot\log\langle\varphi_{t},1\otimes I_{k}\rangle.$
@@
Finally, part (b) follows from Definition 9.4, applied when the sets $\mathcal{X}(\pi_{n},O)$ are empty for some sufficiently small neighbourhood $O$. ∎
@@
Now we can see how the function $\mathrm{h}_{\boldsymbol{\pi}}$ determines strong-quotient limits.
@@
If $\overline{\mathrm{tr}}_{d_{n}}\circ\pi\to\tau$ and $\boldsymbol{\pi}$ strong-quotient converges, then
$\lim_{n}\overline{\Sigma_{k}(\pi_{n})}$ =$=\{\varphi\in\Sigma_{k}(\mathfrak{A}):\ \mathrm{h}_{\boldsymbol{\pi}}(\varphi_{t})\geq 0\ \forall t\geq 0\}$
@@
for each positive integer $k$. ∎
Notions of entropy are often given by concave functions, so it is natural to ask whether the restrictions $\mathrm{h}_{\boldsymbol{\pi}}|\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ have this property. In a restricted sense, they do. First, let us assume again that that $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\to\tau$ and that $\boldsymbol{\pi}$ strong-quotient converges; if these conditions fail then we can always pass to a subsequence that satisfies them. Under these assumptions, we obtain the following from Theorem C together with the log-concavity inequality (3.4) (a consequence of Corollary 3.11).
@@
Let $\varphi,\psi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, and assume moreover that
$t\varphi+(1-t)\psi\ \hbox{is asymptotically associated to}\ \boldsymbol{\pi}\ \hbox{for every}\ t\in[0,1].$ (9.10)
Then $\mathrm{h}_{\boldsymbol{\pi}}$ is concave along the line segment from $\varphi$ to $\psi$:
@@
However, the function $\mathrm{h}_{\boldsymbol{\pi}}$ need not be globally concave on $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, or even on its subset $\Sigma_{k}(\mathfrak{A})$. The problem is that the set $\lim_{n}\overline{\Sigma_{k}(\pi_{n})}$ itself need not be convex. This is analogous to the reason sofic entropy may fail to be additive for Cartesian products in ergodic theory [Aus16]. We illustrate this by an extension of Example 8.2.
@@
Let $\Gamma$ be a countable group with left regular representation $\lambda$. Let $\pi$ be a finite-dimensional irreducible representation of $\Gamma$ that (i) has dimension at least $2$, and (ii) is not approximately contained in $\lambda$. Separately, let $(\rho_{n})_{n\geq 1}$ be an AP sequence for $\Gamma$ that strong-quotient converges to $\lambda$. Suitable examples include randomly generated unitary or permutation representations of free groups, as proved in [CM14] and [BC19], respectively; see also [Mag25].
@@
This is closed for every $k$ and $n$ because it is a continuous image of a finite-dimensional sphere. One can now show that $(\pi_{n})_{n\geq 1}$ strong-quotient converges to $\pi\oplus\lambda$, and so in particular
@@
in the Vietoris topology. However, since $\pi$ is irreducible and not approximately associated to $\lambda$, the same reasoning as for Example 8.2 shows that $\lim_{n}\Sigma_{k}(\pi_{n})$ does not contain averages of the form $(\varphi+\psi)/2$ when $\varphi$ and $\psi$ are both associated to $\pi$ and linearly independent, and so $\lim_{n}\Sigma_{k}(\pi_{n})$ is not convex. Finally, it follows from Corollary 9.20(a) that $\mathrm{h}_{\boldsymbol{\pi}}(\varphi+\tau)$ and $\mathrm{h}_{\boldsymbol{\pi}}(\psi+\tau)$ are both non-negative, but we have
$\mathrm{h}_{\boldsymbol{\pi}}((\varphi+\tau)/2+(\psi+\tau)/2)=-\infty.$
As in Example 9.24, one can show in general that the parts of $\lim_{n}\Sigma_{k}(\pi_{n})$ that lie outside the lower bound on this set provided by Corollary 9.14 are responsible for the failure of convexity.
Nevertheless, if $(\pi_{n})_{n\geq 1}$ strong-quotient converges, then $\lim_{n}\Sigma_{k}(\pi_{n})$ is at least star-shaped around the element $\tau\otimes I_{k}$ (or any other $\lambda$-normal elements of $\Sigma_{k}(\mathfrak{A})$). This is a consequence of Corollary 9.18(a), because $\tau\otimes I_{k}$ does not contribute to the singular part in a convex combination.
Since $\mathrm{h}_{\boldsymbol{\pi}}$ may not be concave on the whole of $\Sigma_{k}(\mathfrak{A})$, it may not be exactly recoverable from its Legendre transform (see [Sim93, Section I.6]). Nevertheless, it might be interesting to investigate how various features of the sequence $(\pi_{n})_{n\geq 1}$ are reflected by that transform, and how it compares with the Legendre transform of $\log\Delta\varphi_{\mathrm{ac}}$ itself. In the analogous setting of modes of convergence for graphs of uniformly bounded degree (see Remarks 8.16 and 9.19), convergence of these Legendre transforms would correspond to ‘right convergence’ [BR11, BCKL13].

### v3

@@ 9.5 Proof of Theorem C
Let $\tau$ be associated to the representation $\lambda$ by the cyclic tracial vector $\xi$, and let $\widetilde{\tau}$ be the normal tracial state on $\lambda(\mathfrak{A})^{\prime}$ defined by $\xi$ as in (3.5). We use the letter $\Delta$ for both (i) the Fuglede–Kadison determinant on $\mathfrak{A}$ defined by $\tau$ and also (ii) the Fuglede–Kadison determinant on log-integrable operators affiliated to $\lambda(\mathfrak{A})^{\prime}$ defined by $\widetilde{\tau}$.
@@
for every $n$. Since $\varphi$ is asymptotically associated to $\boldsymbol{\pi}$ by assumption, so is $\varphi_{\mathrm{sing}}$ by Corollary 9.16. Therefore $\mathcal{X}(\pi_{n},U)$ is nonempty along an infinite subsequence of values of $n$, say $n_{1}<n_{2}<\dots$. Let $\boldsymbol{\pi}^{\prime}$ be the corresponding AP subsequence of $\boldsymbol{\pi}$. For each $n_{i}$, the set $\mathcal{X}(\pi_{n_{i}},O)$ contains a translate of $\mathcal{X}(\pi_{n_{i}},W)$, and so
@@
This lower bound is at least $\log\Delta\varphi_{\mathrm{ac}}$ by applying Step 1 along the AP sequence $\boldsymbol{\pi}^{\prime}$. Since $O$ is arbitrary, this proves that $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ is also at least $\log\Delta\varphi_{\mathrm{ac}}$.
@@ 9.6 Some consequences of Theorem C
  2. b.
    $\pi_{\varphi}\gtrsim\pi_{\varphi_{\mathrm{ac}}}\simeq\lambda^{\oplus k}.$
@@
Part (a) follows directly from Corollary 9.16 and Theorem C.
If $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)>-\infty$, then Theorem C tells us that $\Delta\varphi_{\mathrm{ac}}>0$. Expressing this Fuglede–Kadison determinant in terms of the affiliated operator $T$ from Proposition 4.10, it follows that $\Delta T$ is also positive, and hence $T$ is nonsingular. Since $T$ is self-adjoint, it therefore also has dense image. Since $\pi_{\varphi_{\mathrm{ac}}}$ is equivalent to the subrepresentation of $\lambda^{\oplus k}$ defined by $\overline{\mathrm{img}\,T}$, this proves part (b).
Finally, if $\varphi\geq\psi$ and $\varphi$ is asymptotically associated to $\boldsymbol{\pi}$, then so is $\psi$, and Theorem C shows that part (c) is implied by the inequality $\Delta\varphi_{\mathrm{ac}}\geq\Delta\psi_{\mathrm{ac}}$ from Corollary 3.11(a). ∎
Under the conditions of Theorem C, AP entropy along $\boldsymbol{\pi}$ is really just the logarithm of a Fuglede–Kadison determinant, provided one knows which positive definite functions are asymptotically associated to $\boldsymbol{\pi}$. By passing to a subsequence, one can always assume that $\boldsymbol{\pi}$ strong-quotient converges, and then Theorem C shows that the AP entropy function $\mathrm{h}_{\boldsymbol{\pi}}$ depends only on the values of the limits
@@
In ergodic theory, an important open problem asks whether the sofic entropy of a measure-preserving system along any sofic approximation must always equal either one particular value or $\infty$- ∞. This would mean that the sofic entropy of a measure-preserving system is ‘unambiguous’ in that it does not depend on the choice of sofic approximation provided it is finite. Corollary 9.19 answers the analogous question positively for AP entropy. On the other hand, it is known that the answer is negative for the topological variant of sofic entropy [2]. If the answer is negative for sofic entropy itself, then any general method for evaluating or estimating sofic entropy must depend on the particular sofic approximation through more than just its set of asymptotically associated stationary processes.
@@
Assume again that $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\to\tau$. We can reverse the discussion above by asking whether the function $\mathrm{h}_{\boldsymbol{\pi}}$ determines the topological upper limit of the sequence $(\Sigma_{k}(\pi_{n}))_{n\geq 1}$ for each $k$. This is slightly subtle, because for $\varphi\in\Sigma_{k}(\mathfrak{A})$ the value $\log\Delta_{\tau}\varphi$ may equal $\infty$- ∞ even if $\varphi$ is $\lambda$-normal. We can evade this issue by instead using $\varphi$ to form the perturbations
@@
    If $\varphi_{t}$ is asymptotically associated to $\boldsymbol{\pi}$ for some $t>0$, then this holds for all $t>0$, and we have
@@
    If $\varphi_{t}$ is not asymptotically associated to $\boldsymbol{\pi}$ for any $t>0$, then $\mathrm{h}_{\boldsymbol{\pi}}(\varphi_{t})=-\infty$ for every $t>0$.
@@
for every $t\geq 0$. Therefore, if $\varphi_{t}$ is asymptotically associated to $\boldsymbol{\pi}$ for some $t>0$, then so is $\varphi_{\mathrm{sing}}$, and then so is $\varphi_{s}$ for every other $s$, by two applications of Corollary 9.16.
Now the lower bound on $\mathrm{h}_{\boldsymbol{\pi}}(\varphi_{t})$ in part (a) follows from Corollary 9.19(c) and Proposition 9.17, because $\varphi_{t}\geq\tau\otimes I_{k}$ in the positive definite ordering. On the other hand, since $\Delta_{\tau\otimes\overline{\mathrm{tr}}_{k}}(1\otimes I_{k})=1$, the element $1\otimes I_{k}$ of $\mathrm{M}_{k}(\mathfrak{A})$ is allowed inside the infimum of the variational principle from Proposition 4.13. Combined with Theorem C, this gives
$\mathrm{h}_{\boldsymbol{\pi}}(\varphi_{t})=\log\Delta_{\tau}(\varphi_{t})_{\mathrm{ac}}\leq k\cdot\log\langle\varphi_{t},1\otimes I_{k}\rangle.$
@@
Finally, part (b) follows from Definition 9.4, applied when the sets $\mathcal{X}(\pi_{n},O)$ are empty for some sufficiently small neighbourhood $O$ of $\varphi_{t}$. ∎
@@
Now we can see how the function $\mathrm{h}_{\boldsymbol{\pi}}$ determines topological upper limits.
@@
Assume that $\overline{\mathrm{tr}}_{d_{n}}\circ\pi\to\tau$. Then
$\mathrm{T}\limsup_{n\to\infty}\Sigma_{k}(\pi_{n})$ =$=\{\varphi\in\Sigma_{k}(\mathfrak{A}):\ \mathrm{h}_{\boldsymbol{\pi}}(\varphi_{t})\geq 0\ \forall t\geq 0\}$
@@
for each positive integer $k$, and so $\boldsymbol{\pi}$ strong-quotient converges if and only if we have $\mathrm{h}_{\boldsymbol{\pi}^{\prime}}=\mathrm{h}_{\boldsymbol{\pi}}$ whenever $\boldsymbol{\pi}^{\prime}$ is a subsequence of $\boldsymbol{\pi}$. ∎
In ergodic theory, Kolmogorov–Sinai entropy for single transformations is additive under Cartesian products [157, Theorem 4.23]. By contrast, sofic entropy is subadditive for joinings, but it may fail to be additive, even for a product joining [11]. The same phenomena hold for AP entropy, for essentially the same reasons. Let us assume again that that $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\to\tau$ and that $\boldsymbol{\pi}$ strong-quotient converges; if these conditions fail then we can always pass to a subsequence that satisfies them.
@@
Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, let $\psi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{\ell})_{+}$, and let $\theta$ be a joining of them (recall Definition 4.4). Then
$\mathrm{h}_{\boldsymbol{\pi}}(\theta)\leq\mathrm{h}_{\boldsymbol{\pi}}(\varphi)+\mathrm{h}_{\boldsymbol{\pi}}(\psi).$ (9.10)
If $\theta_{\mathrm{ac}}=\mathrm{diag}(\varphi_{\mathrm{ac}},\psi_{\mathrm{ac}})$, then equality holds provided $\theta$ is asymptotically associated to $\boldsymbol{\pi}$.
Proof.
If $\theta$ is not asymptotically associated to $\boldsymbol{\pi}$, then $\mathrm{h}_{\boldsymbol{\pi}}(\theta)=-\infty$ and there is nothing else to prove. So assume that $\theta$ is asymptotically associated to $\boldsymbol{\pi}$. Both $\varphi$ and $\psi$ are associated to $\pi_{\theta}$, so these are also asymptotically associated to $\boldsymbol{\pi}$. Now the AP entropies of all three maps are given by Theorem C, and the desired conclusions follow from subadditivity and the case of equality for $\log\Delta$: see Corollary 4.14. ∎
Remark 9.25.
Alternatively, the subadditivity of $\mathrm{h}_{\boldsymbol{\pi}}$ may be proved by observing that, for any neighbourhoods $U$ of $\varphi$ and $V$ of $\psi$, there is a neighbourhood $O$ of $\theta$ such that
$\mathcal{X}(\pi,O)\subset\mathcal{X}(\pi,U)\times\mathcal{X}(\pi,V)$
[…]
It is not enough to assume that $\varphi$ and $\psi$ are both separately asymptotically associated to $\boldsymbol{\pi}$ in the last part of Corollary 9.24; see Example 9.27 below.
Notions of entropy are often given by concave functions, so it is natural to ask whether the restrictions $\mathrm{h}_{\boldsymbol{\pi}}|\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ have this property. In a restricted sense, they do. We obtain the following from Theorem C together with the log-concavity inequality (3.4) (a consequence of Corollary 3.11).
Corollary 9.26.
Let $\varphi,\psi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, and assume moreover that ${t\varphi+(1-t)\psi}$ is asymptotically associated to $\boldsymbol{\pi}$ for every $t\in[0,1]$. Then $\mathrm{h}_{\boldsymbol{\pi}}$ is concave along the line segment from $\varphi$ to $\psi$:
@@
However, the function $\mathrm{h}_{\boldsymbol{\pi}}$ need not be globally concave on $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, or even on its subset $\Sigma_{k}(\mathfrak{A})$. The problem is that the set $\lim_{n}\overline{\Sigma_{k}(\pi_{n})}$ itself need not be convex. We illustrate this by an extension of Example 8.2.
@@
Let $\Gamma$ be a countable group with left regular representation $\lambda$. Let $\pi$ be a finite-dimensional irreducible representation of $\Gamma$ that (i) has dimension at least $2$, and (ii) is not approximately contained in $\lambda$. Separately, let $(\rho_{n})_{n\geq 1}$ be an AP sequence for $\Gamma$ that strong-quotient converges to $\lambda$. Suitable examples include randomly generated unitary or permutation representations of free groups, as proved in [35] and [21], respectively; see also [113].
@@
This is closed for every $k$ and $n$ because it is a continuous image of a finite-dimensional sphere. One can now check that $(\pi_{n})_{n\geq 1}$ strong-quotient converges to $\pi\oplus\lambda$, and so in particular
@@
in the Vietoris topology. However, since $\pi$ is irreducible and not approximately associated to $\lambda$, it follows that $\lim_{n}\Sigma_{2}(\pi_{n})$ does not contain $\mathrm{diag}(\varphi,\psi)$ when $\varphi$ and $\psi$ are linearly independent elements of $\Sigma_{1}(\pi)$. Moreover, by the same reasoning as for Example 8.2, it also follows that $\lim_{n}\Sigma_{1}(\pi_{n})$ does not contain $(\varphi+\psi)/2$, so $\lim_{n}\Sigma_{1}(\pi_{n})$ is not convex. Finally, Corollary 9.21(a) gives that $\mathrm{h}_{\boldsymbol{\pi}}(\varphi+\tau)$ and $\mathrm{h}_{\boldsymbol{\pi}}(\psi+\tau)$ are both non-negative, but we have
$\mathrm{h}_{\boldsymbol{\pi}}(\mathrm{diag}(\varphi+\tau,\psi+\tau))=\mathrm{h}_{\boldsymbol{\pi}}((\varphi+\tau)/2+(\psi+\tau)/2)=-\infty.$
As suggested by Example 9.27, one can show in general that the parts of $\lim_{n}\Sigma_{k}(\pi_{n})$ that lie outside the lower bound on this set provided by Corollary 9.15 are responsible for the failure of convexity.
Nevertheless, if $(\pi_{n})_{n\geq 1}$ strong-quotient converges, then $\lim_{n}\Sigma_{k}(\pi_{n})$ is at least star-shaped around the element $\tau\otimes I_{k}$ (or any other $\lambda$-normal elements of $\Sigma_{k}(\mathfrak{A})$). This is a consequence of Corollary 9.16, because $\tau\otimes I_{k}$ does not contribute to the singular part in a convex combination.
Since $\mathrm{h}_{\boldsymbol{\pi}}$ may not be concave on the whole of $\Sigma_{k}(\mathfrak{A})$, it may not be exactly recoverable from its Legendre transform. Nevertheless, it might be interesting to investigate how various features of the sequence $(\pi_{n})_{n\geq 1}$ are reflected by that transform, and how it compares with the Legendre transform of $\log\Delta\varphi_{\mathrm{ac}}$ itself. See [135, Section I.6] for a general account of infinite-dimensional Legendre transforms with a view towards statistical mechanics, or [34] for the resulting variational principle in the case of sofic entropy. In the analogous setting of modes of convergence for graphs of uniformly bounded degree (see Remarks 8.13 and 9.20), convergence of these Legendre transforms would correspond to ‘right convergence’ [20, 23].